# Phases and phase transition in Grover’s algorithm with systematic noise

### v1

@@ Abstract
While limitations on quantum computation by Markovian environmental noise are well-understood in generality, their behavior for different quantum circuits and noise realizations can be less universal. Here we consider a canonical quantum algorithm - Grover’s algorithm for unordered search on $L$ qubits - in the presence of systematic noise. This allows us to write the behavior as a random Floquet unitary, which we show is well-characterized by random matrix theory (RMT). The RMT analysis enables analytical predictions for phases and phase transitions of the many-body dynamics. We find two separate transitions. At moderate disorder $\delta_{c,\mathrm{gap}}\sim L^{-1}$, there is a ergodicity breaking transition such that a finite-dimensional manifold remains non-ergodic for $\delta<\delta_{c,\mathrm{gap}}$. Computational power is lost at a much smaller disorder, $\delta_{c,\mathrm{comp}}\sim L^{-1/2}2^{-L/2}$. We comment on relevance to non-systematic noise in realistic quantum computers, including cold atom, trapped ion, and superconducting platforms.
@@ I Introduction
Until such a time as quantum error correction is achieved, it remains important to understand the role of noise in creating issues for known quantum algorithms. One such algorithm is Grover’s algorithm for unstructured database search [4]. It has been the subject of intense study and has, in fact, been implemented on various NISQ devices [5, 6, 7]. The effects of noise on Grover’s search has been studied using different models and efforts have been made towards putting limits on the performance of the algorithm. For example, Long et al. [8] determined that the size of the database is constrained by $\mathcal{O}(1/\delta^{2})$ where $\delta$ is an error parameter representing gate imperfection. In the presence of random phase error in the oracle gate, Shenvi et al. [9] have shown that in order to maintain the probability of success, the error must decrease as $k^{1/4}$ when the database size is increased by a factor of $k$. In case of depolarizing noise channel, Salas [10] found that the noise threshold behaves as $N^{-1.1}$, where $N$ is the size of the database. Norman and Ataba [11] determined that the allowed noise rate scales as $N^{-2/3}$ if Gaussian noise is added at each step of the algorithm. Finally, Shapira et al. [12] showed that in case of unitary noise characterized by standard deviation $\epsilon$, must satisfy $\epsilon<\mathcal{O}(L^{-1/2}N^{-1/4})$ where $N=2^{L}$, in order for the algorithm to maintain a significant efficiency.
@@
These cases primarily studied aphysical noise models on the long-range many-qubit Grover gates and consistently found that exponentially small error rates prevent the algorithm from succeeding. These error models make generalizing the results rather challenging, leaving open the question whether other physically realistic error models might have improved performance bounds. Furthermore, the standard Grover algorithm involves a call to an oracle which is a non-$k$-local operation that is aphysical in most realizations. We will consider the impact of errors on a circuit in which the oracle is unraveled into $k$-local unitary gates (cf. Figure 2), which comes at the cost of larger depth.
@@
In order to get a better analytical handle on the error-tolerance of Grover’s algorithm, we consider here the effects of systematic gate noise, i.e., undesired noise in the quantum gates that is repeated systematically throughout the course of the algorithm. Systematic noise naturally occurs in most experiments due to calibration issues. For our purposes, it imbues the noisy algorithm with a time-periodic (Floquet) structure, such that we can find exact scaling of target state dynamics, noise thresholds, and other physically relevant quantities by employing both analytical and numerical approaches. We show that these signatures can be captured by a simple random-matrix Hamiltonian whose perturbative effect on the Floquet spectrum describes gap-closing transitions that determine the response to noise. We find that for a database of size $N=2^{L}$, the algorithm loses its computational power at a noise strength that scales as $L^{-1/2}N^{-1/2}$. Additionally, an ergodicity transition occurs at a noise strength that scales as $1/L$ with the system size. Using state-of-the-art numerics, we show that these analytical predictions align perfectly with the exact numerical solution and comment on the relevance to using Grover’s algorithm – as well as related amplitude amplification techniques – in the NISQ era.
@@ II Model
We start by reviewing the basic formulation of Grover’s algorithm. Suppose we want to search for an item $w$ in an unstructured database with $N=2^{L}$ items. For simplicity, let us assume that the target item $w$ is unique. The search problem can be expressed by using an indicator function $f(x)$ that takes integer inputs $x$, with $10\leq x\leq N-$ 1, such that
$f(x)=\begin{cases}1&\text{if $x$ is a solution, i.e., $x=w$}\\ 0&\text{otherwise.}\end{cases}$ (1)
In case of Grover’s algorithm, this indicator function is implemented using a unitary operator $U_{w}$, known as the oracle, whose action on the computational basis states is
$|x\rangle\xrightarrow{U_{w}}(-1)^{f(x)}|x\rangle.$ (2)
The complexity of an algorithm can be measured by the number of times the function $f(x)$ is evaluated. Classical algorithms require $\sim N/2$ “time steps” – calls to a circuit that evaluates the function $f(x)$ – on average to find the target state. Grover’s algorithm requires $\mathcal{O}(\sqrt{N})$ evaluations of the oracle $O$ to obtain the target state, and thus provides quadratic advantage over the classical counterparts.
Instead of the elements of the database directly, Grover’s algorithm focuses on the indices of the elements, such that the database can be represented using $L$ qubits as $\{|0\rangle,|1\rangle,\cdots,|2^{L}-1\rangle\}$. The algorithm starts with the uniform superposition state $|x\rangle=\frac{1}{\sqrt{N}}\sum^{N-1}_{j=0}|j\rangle$, and then amplifies the amplitude of the target state $|w\rangle$ using the action of the two unitary operators. First it applies the oracle, which flips the sign of the state $|w\rangle$,
@@
The combined action of these two reflections are written in terms of the Grover operator as $G_{w}=U_{x}U_{w}$, which gradually rotates the initial state $|x\rangle$ towards the target state $|w\rangle$. The action of $G_{w}$ on the state of the system can be interpreted geometrically by writing the state of the system at $t=0$ as
@@
In other words, $G_{w}$ rotates the wave function in the two dimensional subspace spanned by $|\bar{x}_{w}\rangle$ and $|w\rangle$ by angle $2\theta$. From Eq. (7), we get the probability of measuring the target $|w\rangle$ after $t$ steps as,
@@ II.1 Floquet spectrum
The quasi-energies $\phi_{\alpha}$ are defined modulo $2\pi/T$, where period $T\equiv 1$ represents the time required to apply $G_{w}$ on the qubits.
Before we write down the exact eigenstates and quasi-energies for the Grover operator, we state a useful theorem that simplifies our analysis: the spectrum of the Grover operator is independent of the target state $w$. To prove this, consider conjugation by the unitary matrix
@@
where the notation $j\in w$ indicates to take a product over Pauli $X$ operators only acting on sites where $w$ is in state $1$. Then
@@
Since $G_{0}$ and $G_{w}$ are unitarily equivalent, their eigenvalues are identical. We will therefore focus on target state $w=0$ (state $|0\rangle$ on all qubits) for the remainder of the work.
The Grover operator $G_{0}$ has two eigenstates with non-zero quasi-energies, which we will refer to as the special states,
$G_{0}\frac{1}{\sqrt{2}}(|0\rangle\pm i|\bar{x}\rangle)$ $=$ $[\frac{2}{N}-1\mp i\frac{2\sqrt{N-1}}{N}]\frac{1}{\sqrt{2}}(|0\rangle\pm i|\bar{x}\rangle)$ (13)
$=$ $e^{-i\phi_{F\pm}}\frac{1}{\sqrt{2}}(|0\rangle\pm i|\bar{x}\rangle),$
@@
$\phi_{F_{\pm}}=-\pi\mp\arctan\left(\frac{2\sqrt{N-1}}{N-2}\right)=-\pi\mp 2\theta,$ (14)
@@
$|\bar{x}\rangle\equiv|\bar{x}(w=0)\rangle=\frac{1}{\sqrt{N-1}}\sum^{N-1}_{j=1}|j\rangle.$ (15)
@@
Noting that $|\bar{x}\rangle$ is a “zero-momentum” superposition of the non-target states, we choose the following basis for the degenerate bulk subspace:
@@
where $k_{n}=(2\pi n)/(N-1)$, and $n=1,2,\cdots,N-2$. In the absence of noise, these states are gapped out and decoupled from the dynamics.
@@ II.2 Gate decomposition
Figure 1: Circuit for the four qubit Grover operator $G_{0}$.
In Figure 1, the quantum circuit for Grover’s operator $G_{0}$ for four qubits is shown. The oracle and the Grover’s diffusion operators are written in terms of a multi-controlled $X$, a.k.a. $L-$qubit Toffoli, gate for ease of decomposition, as such gates are more commonly studied in the literature than the desired multi-controlled $Z$ gate. Decomposing the $L$–qubit Toffoli can be done many different ways, such as using automated transpilation methods (e.g. ‘transpile’ in Qiskit) or via the general algorithm described in the classic paper by Barenco et al. [19]. However, these decompositions result in a circuits whose depth scales non-optimally with system size; it appears to scale exponentially with $L$ using Qiskit transpilation and is known to scale as $L^{2}$ using the Barenco algorithm. In anticipation of improving performance on the noisy algorithm, we instead use the linear depth decomposition algorithm described in [20]. For a multi-controlled $X$ gate with $L$ qubits, this algorithm provides a decomposition with circuit depth $208L-$ 20 and with $52L^{2}-6L+$ 5 controlled $x$-rotation gates, hereafter denoted by $CR_{x}(\theta)$. An example for the case of $L=4$ is shown in Figure 2. Note that, while the depth is linear, it relies on the ability to perform multiple finite-range  [^1] The decomposition shown requires gate range proportional to $L$, but the authors argue that finite-range gates still allow depth $O(L)$. control gates in parallel that share the same control qubit. If these gates cannot be parallelized, the results below will be unchanged, but any realistic quantum computer that includes non-unitary noise ($T_{1}$, $T_{2}$, etc.) will have its performance degraded. Note also that here a single Floquet unitary already has extensive depth $\sim L$, as opposed to more conventional Floquet operators whose depth is usually $L$-independent.
Figure 2: Example decomposition of the four-qubit-controlled $X$ gate in Figure 1 in linear depth using the algorithm described in [20]. The gates with same control qubit are shown together.
@@ II.3 Systematic noise
We consider a systematic noise model with random under or over rotation of each single or two-qubit quantum gate,
@@
where the index $k$ refers to the $k^{th}$ gate in the decomposition of the Grover operator, $\epsilon_{k}$ is drawn from the uniform distribution $[-1,1]$ and $\delta$ is the noise strength. Importantly, each Floquet cycle has an identical noise configuration, so that the operator $G_{w}$ is identical across time steps. This ensures that each iteration of the Grover algorithm is a (random) Floquet problem.
@@
represents a rotation by angle $\theta$ about the $\sigma_{k}$ axis. Then we define the noisy gates as a random over/under-rotation,
@@ III.1 Spectrum of noisy Grover operator
To begin exploring the effect of system noise on the performance of the algorithm, we numerically calculate the Floquet quasi-energies and eigenstates of the Grover operator in presence of systematic noise. To probe thermalization of the bulk states, we calculate the bipartite von Neumann entanglement entropy $S_{A}$ of each eigenstate. We choose a subsystem $A$ of $L/2$ consecutive qubits (with periodic boundary conditions) and calculate the entropy $S_{A}=-\text{Tr}(\rho_{A}\ln\rho_{A})$, where $\rho_{A}=\text{Tr}_{\bar{A}}(|\psi\rangle\langle\psi|)$ is the reduced density matrix in eigenstate $|\psi\rangle$. For each eigenstate, the entanglement entropy is averaged over the $L/2$ independent choices of subsystem $A$
@@
Figure 3: (a) Quasienergies $(\phi_{F})$ of the Grover operator as a function of noise strength $(\delta)$ for $L=12$. The colorbar represents the average bipartite von Neumann entanglement entropy $S_{A}$ normalized by its Page value [22], $S_{\rm Page}=3.66$ for $L=12$. The data is shown for a single realization of systematic noise $\{\epsilon_{k}\}$, which is kept constant for different $\delta$ values. The two critical noise values $\delta_{c,\mathrm{comp}}$ and $\delta_{c,\mathrm{gap}}$ are shown with vertical lines.(b) Entanglement entropy of the bulk quasienergies at noise strength $\delta=0.08$.
@@
An example of the quasi-energies and their entanglement entropies are shown for a single noise realization in Figure 3. The bulk quasi-energies have their degeneracy broken by the noise. Moreover, these bulk states show clear linear dependence on the noise strength $\delta$, which comes from the exponential degeneracy of the unperturbed spectrum. This suggests that the bulk properties are well-described by first-order perturbation theory, even at relatively large $\delta$, which we use in the next section to derive an effective Hamiltonian for the quasi-energy spectrum. Because the bulk spectrum remains far from the edge of the Floquet Brillouin zone at $\pi$ quasienergy, the effective Hamiltonian is time-independent despite coming from a Floquet problem.
@@
In addition to this linear $\delta$-dependence of quasi-energy, the bulk eigenstates show entanglement characteristic of a thermal quantum system, where they should match the thermal entropy $S(E)$. The entanglement peaks in the middle of the bulk spectrum at approximately its maximum value $S_{\mathrm{Page}}=\left[L\ln(2)-1\right]/2$ [22]. As expected, the entropy density approaches zero at either edge of the bulk spectrum. Further evidence for thermalizing (random matrix) behavior will be given in Section III.2.1.
@@
Finally, below some critical value of noise strength $\delta<\delta_{c,\rm gap}$, the special states remain gapped from the bulk. Since Grover’s algorithm involves coherent oscillations within the special state manifold, this suggests that there is a “computing” phase at small $\delta$ where finite probability of finding the target state will survive the systematic noise. We will indeed show that this is true at small enough $\delta$ though, as we will see in our analysis of the special states (Section III.4), the computing transition happens at a much smaller critical value $\delta_{c,\rm comp}\ll\delta_{c,\rm gap}$. Anticipating the results that we are about to show, $\delta_{c,\rm gap}$ will in fact correspond to an ergodicity transition; for $\delta<\delta_{c,\mathrm{gap}}$, initial states within the special state manifold do not thermalize at late time, while for $\delta>\delta_{c,\mathrm{gap}}$, all states thermalize to infinite temperature.

### v2

@@ Abstract
While limitations on quantum computation under Markovian environmental noise are well-understood in generality, their behavior for different quantum circuits and noise realizations can be less universal. Here we consider a canonical quantum algorithm - Grover’s algorithm for unordered search on $L$ qubits - in the presence of systematic noise. This allows us to write the evolution as a random Floquet unitary, which we show is well-characterized by random matrix theory (RMT). The RMT analysis enables analytical predictions for different regimes and critical disorders of the many-body dynamics. We find two critical disorders. At small disorder, finite-dimensional manifold remains nonergodic as long as the noise is smaller than an ergodicity breaking transition, which scales as $O(L^{-1})$. Computational power is lost at a much smaller disorder that scales as $O(2^{-L/2})$. We comment on relevance to non-systematic noise in realistic quantum computers, including cold atom, trapped ion, and superconducting platforms.
@@ I Introduction
Until such a time as quantum error correction is achieved, it remains important to understand the role of noise in creating issues for known quantum algorithms. One such algorithm is Grover’s algorithm for unstructured database search [4]. It has been the subject of intense study and has, in fact, been implemented on various NISQ devices [5, 6, 7]. The effects of noise on Grover’s search have been studied using different models and efforts have been made towards putting limits on the performance of the algorithm. For example, Long et al[8] determined that the size of the database is constrained by $\mathcal{O}(1/\delta^{2})$ where $\delta$ is an error parameter representing gate imperfection. In the presence of random phase error in the oracle gate, Shenvi et al[9] have shown that in order to maintain the probability of success, the error must decrease as $k^{1/4}$ when the database size is increased by a factor of $k$. In the case of depolarizing noise channel, Salas [10] found that the noise threshold behaves as $N^{-1.1}$, where $N$ is the size of the database. Norman and Ataba [11] determined that the allowed noise rate scales as $N^{-2/3}$ if Gaussian noise is added at each step of the algorithm. Finally, Shapira et al[12] showed that for Grover’s algorithm with $L$ qubits, with unitary noise characterized by standard deviation $\epsilon$, must satisfy $\epsilon<\mathcal{O}(L^{-1/2}N^{-1/4})$ where $N=2^{L}$, in order for the algorithm to maintain a significant efficiency. It was also shown that any NISQ algorithm will fail to achieve the quadratic speedup over classical algorithms in the unstructured database search problem[13].
@@
These cases primarily studied unphysical noise models on the long-range many-qubit Grover gates and consistently found that exponentially small error rates prevent the algorithm from succeeding. These error models make generalizing the results rather challenging, leaving open the question of whether other physically realistic error models might have improved performance bounds. Furthermore, the standard Grover algorithm involves a call to an oracle which is a non-$k$-local operation that is unphysical in most realizations. In practice, such non-$k-$local gates are implemented by decomposing it in terms of $k-$local gates ($k=1,2$ in most cases that act on one qubit, or two qubits that are not necessarily adjacent to each other). We will consider the impact of errors on a circuit in which the oracle is unraveled into single qubit gates and $2$-local unitary gates (cf. Fig 2), which comes at the cost of larger depth.
@@
In order to get a better analytical handle on the error-tolerance of Grover’s algorithm, we consider here the effects of systematic gate noise, i.e., undesired noise in the quantum gates that is repeated systematically throughout the course of the algorithm. Systematic noise naturally occurs in most experiments due to calibration issues. For our purposes, it imbues the noisy algorithm with a time-periodic (Floquet) structure, such that we can find the exact scaling of target state dynamics, noise thresholds, and other physically relevant quantities by employing both analytical and numerical approaches. We show that these signatures can be captured by a simple random-matrix Hamiltonian whose perturbative effect on the Floquet spectrum describes gap-closing transitions that determine the response to noise. We find that for a database of size $N=2^{L}$, the algorithm loses its computational power at a noise strength that scales as $\mathcal{O}(N^{-1/2})$. Additionally, an ergodicity transition occurs at a noise strength that scales as $\mathcal{O}(L^{-1})$ with the system size. We show that these analytical predictions align perfectly with the exact numerical solution and comment on the relevance to using Grover’s algorithm – as well as related amplitude amplification techniques – in the NISQ era.
@@ II Model
We start by reviewing the basic formulation of Grover’s algorithm. Suppose we want to search for an item $w$ in an unstructured database with $N=2^{L}$ items. For simplicity, let us assume that the target item $w$ is unique. The search problem can be expressed by using an indicator function $f(u)$ that takes integer inputs $u$, with $10\leq u\leq N-$ 1, such that
$f(u)=\begin{cases}1&\text{if $u$ is the solution, i.e., $u=w$}\\ 0&\text{otherwise.}\end{cases}$ (1)
In the case of Grover’s algorithm, this indicator function is implemented using a unitary operator $U_{w}$, known as the oracle, whose action on the computational basis states is
$|u\rangle\xrightarrow{U_{w}}(-1)^{f(u)}|u\rangle.$ (2)
The complexity of an algorithm can be measured by the number of times the function $f(u)$ is evaluated. Classical algorithms require $\sim N/2$ “time steps” – calls to a circuit that evaluates the function $f(u)$ – on average to find the target state. Grover’s algorithm requires $\mathcal{O}(\sqrt{N})$ evaluations of the oracle $U_{w}$ to obtain the target state, and thus provides quadratic speed up of query complexity over the classical counterparts.
Instead of the elements of the database directly, Grover’s algorithm focuses on the indices of the elements, such that the database can be represented using $L$ qubits as $\{|0\rangle,|1\rangle,\cdots,|2^{L}-1\rangle\}$. The algorithm starts with the uniform superposition state $|x\rangle=\frac{1}{\sqrt{N}}\sum^{N-1}_{j=0}|j\rangle$, and then amplifies the amplitude of the target state $|w\rangle$ using the action of the two unitary operators. First, it applies the oracle, which flips the sign of all the states except for the state $|w\rangle$,
@@
The combined action of these two reflections is written in terms of the Grover operator as $G_{w}=U_{x}U_{w}$, which gradually rotates the initial state $|x\rangle$ towards the target state $|w\rangle$. The action of $G_{w}$ on the state of the system can be interpreted geometrically by writing the state of the system at $t=0$ as
@@
In other words, $G_{w}$ rotates the wave function in the two-dimensional subspace spanned by $|\bar{x}_{w}\rangle$ and $|w\rangle$ by angle $2\theta$. From Eq. (7), we get the probability of measuring the target $|w\rangle$ after $t$ steps as,
@@ II.1 Floquet spectrum
The quasi-energies $\phi_{\alpha}$ are defined modulo $2\pi/T$, where period $T\equiv 1$ represents the time required to apply $G_{w}$ on the qubits. The actual value of time $T$ is irrelevant to this analysis because we can express everything using dimensionless phases rather than working with energies.
Before we write the exact eigenstates and quasi-energies for the Grover operator, we state a useful theorem that simplifies our analysis: the spectrum of the Grover operator is independent of the target state $w$. To prove this, consider conjugation by the unitary matrix
@@
where the notation $j\in w$ indicates to take a product over Pauli $X$ operators acting only on sites where $w$ is in state $1$. Then
@@
Since $G_{0}$ and $G_{w}$ are unitarily equivalent, their eigenvalues are identical. We will therefore focus on target state $|0\rangle^{\otimes L}\equiv|\textbf{0}\rangle$ for the remainder of the work.
The Grover operator $G_{0}$ has two eigenstates with nonzero quasi-energies, which we will refer to as the special states,
$G_{0}\frac{1}{\sqrt{2}}(|\textbf{0}\rangle\pm i|\bar{x}\rangle)$ $=$ $[\frac{2}{N}-1\mp i\frac{2\sqrt{N-1}}{N}]\frac{1}{\sqrt{2}}(|\textbf{0}\rangle\pm i|\bar{x}\rangle)$ (13)
$=$ $e^{-i\phi_{F\pm}}\frac{1}{\sqrt{2}}(|\textbf{0}\rangle\pm i|\bar{x}\rangle),$
@@
$\phi_{F_{\pm}}=\pm(\pi-\arctan\left(\frac{2\sqrt{N-1}}{N-2}\right))=\pm(\pi-2\theta),$ (14)
@@
$|\bar{x}\rangle\equiv|\bar{x}(w=\textbf{0})\rangle=\frac{1}{\sqrt{N-1}}\sum^{N-1}_{j=1}|j\rangle.$ (15)
@@
Noting that $|\bar{x}\rangle$ is a “zero-momentum” superposition of the nontarget states, we choose the following basis for the degenerate bulk subspace:
@@
where $k_{n}=(2\pi n)/(N-1)$, and $n=1,2,\cdots,N-2$. In the absence of noise, these states are gapped out i.e. there is an energy difference (gap) between these bulk states and the states involved in the dynamics, and the evolution of the system takes place in the subspace that does not involve these states.
@@ II.2 Gate decomposition
Figure 1: Circuit for the four qubit Grover operator $G_{0}$. $X,Z$ and $H$ are the Pauli $X$, Pauli $Z$ and the Hadamard gate respectively.
In Figure 1, a quantum circuit for Grover’s operator $G_{0}$ for four qubits is shown. The oracle and the Grover’s diffusion operators are written in terms of a multi-controlled $X$, a.k.a. $L-$qubit Toffoli, gate for ease of decomposition, as such gates are more commonly studied in the literature than the desired multi-controlled $Z$ gate. Decomposing the $L$–qubit Toffoli can be done in many different ways, such as using automated transpilation methods (e.g. ‘transpile’ in Qiskit) or via the general algorithm described in the classic paper by Barenco et al[20]. However, these decompositions result in circuits whose depth scales nonoptimally with system size; it appears to scale exponentially with $L$ using Qiskit transpilation and is known to scale as $L^{2}$ using the Barenco algorithm. In anticipation of improving performance on the noisy algorithm, we instead use the linear depth decomposition algorithm described in [21]. For a multi-controlled $X$ gate with $L$ qubits, this algorithm provides a decomposition with circuit depth $208L-$ 20 and with $52L^{2}-6L+$ 5 controlled $x$-rotation gates [^1] For a rigorous argument for the linear depth of this decomposition, see [21, PhysRevA.106.042602], hereafter denoted by $CR_{x}(\theta)$. An example for the case of $L=4$ is shown in Fig 2. Note that, while the depth is linear, it relies on the ability to perform multiple finite-range  [^2] The decomposition shown requires gate range proportional to $L$, but the authors argue that finite-range gates still allow depth $O(L)$. control gates in parallel. Note also that here a single Floquet unitary already has extensive depth $\sim L$, as opposed to more conventional Floquet operators whose depth is usually $L$-independent.
Figure 2: Example decomposition of the four-qubit-controlled $X$ gate in Fig 1 in linear depth using the algorithm described in [21]. The gates with the same control qubit are shown together.
@@ II.3 Systematic noise
We consider a coherent systematic noise model with random under or over-rotation of each single or two-qubit quantum gate,
@@
where the index $k$ refers to the $k$th gate in the decomposition of the Grover operator, $\epsilon_{k}$ is drawn from the uniform distribution $[-1,1]$ and $\delta$ is the noise strength. Importantly, each Floquet cycle has an identical noise configuration, so that the operator $G_{w}$ is identical across time steps. This ensures that each iteration of the Grover algorithm is a (random) Floquet problem.
@@
represents a rotation by angle $\theta$ about the $\sigma_{k}$ axis. Then we define the noisy gates as a random over -or under-rotation,
@@ III.1 Spectrum of noisy Grover operator
To begin exploring the effect of system noise on the performance of the algorithm, we numerically calculate the Floquet quasi-energies and eigenstates of the Grover operator in the presence of systematic noise. To probe the thermalization of the bulk states, we calculate the bipartite von Neumann entanglement entropy $S_{A}$ of each eigenstate. We choose a subsystem $A$ of $L/2$ consecutive qubits (with periodic boundary conditions) and calculate the entropy $S_{A}=-\text{Tr}(\rho_{A}\ln\rho_{A})$, where $\rho_{A}=\text{Tr}_{\bar{A}}(|\psi\rangle\langle\psi|)$ is the reduced density matrix in eigenstate $|\psi\rangle$. For each eigenstate, the entanglement entropy is averaged over the $L/2$ independent choices of subsystem $A$.
@@
Figure 3: (a) Quasienergies $(\phi_{F})$ of the Grover operator as a function of noise strength $(\delta)$ for $L=12$. The colorbar represents the average bipartite von Neumann entanglement entropy $S_{A}$ normalized by its Page value [24], $S_{\rm Page}=3.66$ for $L=12$. The data are shown for a single realization of systematic noise $\{\epsilon_{k}\}$, which is kept constant for different $\delta$ values. The two critical noise values $\delta_{c,\mathrm{comp}}$ and $\delta_{c,\mathrm{gap}}$ are shown with vertical lines.(b) Entanglement entropy of the bulk quasienergies at noise strength $\delta=0.08$.
@@
An example of the quasi-energies and their entanglement entropies are shown for a single noise realization in Figure 3. The bulk quasi-energies have their degeneracy broken by the noise. Moreover, the quasienergies of these bulk states show a clear linear dependence on the noise strength $\delta$, which comes from the exponential degeneracy of the unperturbed spectrum. This suggests that the bulk properties are well-described by first-order perturbation theory, even at relatively large $\delta$, which we use in the next section to derive an effective Hamiltonian for the quasi-energy spectrum.
@@
In addition to this linear $\delta$-dependence of quasi-energy, the bulk eigenstates show entanglement characteristic of a thermal quantum system, where they should match the thermal entropy $S(E)$. The entanglement peaks in the middle of the bulk spectrum at approximately its maximum value $S_{\mathrm{Page}}=\left[L\ln(2)-1\right]/2$ [24]. As expected, the entropy density approaches zero at either edge of the bulk spectrum. Further evidence for thermalizing (random matrix) behavior will be given in Sec III.2.1.
@@
Finally, below some critical value of noise strength $\delta<\delta_{c,\rm gap}$, the special states remain gapped from the bulk. Since Grover’s algorithm involves coherent oscillations within the special state manifold, this suggests that there is a “computing” regime at small $\delta$ where finite probability of finding the target state will survive the systematic noise. We will indeed show that this is true at small enough $\delta$ though, as we will see in our analysis of the special states (Sec III.4), the computing transition happens at a much smaller critical value $\delta_{c,\rm comp}\ll\delta_{c,\rm gap}$. Anticipating the results that we are about to show, $\delta_{c,\rm gap}$ will in fact correspond to an ergodicity transition; for $\delta<\delta_{c,\mathrm{gap}}$, the initial states within the special states subspace do not reach their thermal equilibrium expectation values i.e. retain memory of their initial conditions even after a long time, while for $\delta>\delta_{c,\mathrm{gap}},$ all states obtain their thermal expectation values.